# MV2Cyl: Reconstructing 3D Extrusion Cylinders from Multi-View Images

### v1

@@ Abstract
We present MV2Cyl, a novel method for reconstructing 3D from 2D multi-view images, not merely as a field or raw geometry but as a sketch-extrude CAD model. Extracting extrusion cylinders from raw 3D geometry has been extensively researched in computer vision, while the processing of 3D data through neural networks has remained a bottleneck. Since 3D scans are generally accompanied by multi-view images, leveraging 2D convolutional neural networks allows these images to be exploited as a rich source for extracting extrusion cylinder information. However, we observe that extracting only the surface information of the extrudes and utilizing it results in suboptimal outcomes due to the challenges in the occlusion and surface segmentation. By synergizing with the extracted base curve information, we achieve the optimal reconstruction result with the best accuracy in 2D sketch and extrude parameter estimation. Our experiments, comparing our method with previous work that takes a raw 3D point cloud as input, demonstrate the effectiveness of our approach by taking advantage of multi-view images.
@@ 2 Related Work
Another line of work tackles fitting 3D parametric curves, where a common strategy is to first detect edges [69] and corners [35, 69, 39] from an input point cloud, and then fit or predict parametric curves from the initial proposals. A recent work, NEF [67], takes multi-view images as input and introduces a self-supervised approach to the parametric curve fitting problem using neural fields, thus allowing them to optimize without needing clean point cloud data. Despite having multi-view images as input, NEF only tackles recovering parametric curves and not extrusion cylinders, which are a basic building block for the sketch-extrude reverse engineering task and are the focus of our work.
@@ 3 MV2Cyl: Reconstructing 3D Extrusion Cylinders from Multi-View Images
We propose a method that reconstructs 3D extrusion cylinders from multi-view images without relying on raw 3D geometry as input. The idea is to leverage 2D neural networks to learn 2D priors that provide 3D extrusion information (Sec. 3.2), i.e. extrusion curves and surfaces. The integration of the information into 3D by optimizing neural fields (Sec. 3.3). MV2Cyl is a combination of a curve field and a surface field that is used to recover the parameters and reconstruct the extrusion cylinders (Sec. 3.4) only given 2D input images.
@@ 3.2 Learning 2D Priors for 3D Extrusions
We observe that detecting and extracting curves on the base planes of each extrusion cylinder segment provides sufficient information to recover the parameters and reconstruct the extrusion cylinders. Moreover, (feature) curves have been a longstanding and well-explored problem dating back to classical computer vision literature [46, 73, 43] thanks to their strong expressiveness. This leads us to learn a 2D curve prior that provides more discriminative and detailed outputs to make reverse engineering more robust.
@@
$\mathcal{L}_{\text{2D}}^{\text{curve}}=\lambda_{CE}\mathcal{L}_{\text{cross-entropy}}+\lambda_{\text{focal}}\mathcal{L}_{\text{focal}}+\lambda_{\text{dice}}\mathcal{L}_{\text{dice}}.\vspace{-0.2cm}$ (2)
Additional details about the segmentation framework are available in the appendix A5.1.
@@ 3.3 Integrating 2D Segments into a 3D Field
Figure 2: Overview of the learned surface and curve fields. (Left-to-Right) Existence field of surface, instance semantic field of surface, start-end semantic field of surface, existence field of curve, instance semantic field of curve, and start-end semantic field of curve.
To establish correspondences and collate the information extracted from multi-view images, a natural approach to integrate 2D information into a consistent 3D representation is through learning a 3D field [40, 60, 67, 44]. For each of our 2D priors $\mathcal{M}^{\text{surface}}$ and $\mathcal{M}^{\text{curve}}$, we learn an existence field $\mathcal{F}$ and an attribute field $\mathcal{A}$ as detailed below, respectively.
Existence Field $\mathcal{F}$
We learn an existence field for surfaces $\mathcal{F}^{\text{surface}}$ and curves $\mathcal{F}^{\text{curve}}$. Given a query point in 3D space $\mathbf{x}\in\mathbb{R}^{3}$, the existence field network $\mathcal{F}:\mathbb{R}^{3}\mapsto\mathbb{R}$ outputs a scalar value that indicates how likely the 3D point $\mathbf{x}$ is on a surface or a curve for $\mathcal{F}^{\text{surface}}$ and curves $\mathcal{F}^{\text{curve}}$, respectively. To optimize the field differentiably with only 2D images, we use volume rendering [40, 5] and a learnable transformation [67] to convert the existence field $\mathcal{F}(\mathbf{x})$ to an opacity field $\sigma(\mathbf{x})$. This transformation is in the form of a learnable sigmoid function given as:
$\sigma(\mathbf{x})=\alpha\cdot\left(1+e^{-\zeta\cdot(\mathcal{F}(\mathbf{x})-\beta)}\right)^{-1}$ (3)
where $\alpha$ is a learnable parameter that adjusts the existence scale, and $\beta$ and $\zeta$ are constant hyperparameters. We use the same transformation in Eq. 3 for the surface field $\mathcal{F}^{\text{surface}}$ and the curve field $\mathcal{F}^{\text{curve}}$, resulting in corresponding opacity fields $\sigma^{\text{surface}}$ and $\sigma^{\text{curve}}$. We then directly render the existence by using the volume rendering equation to get the expected surface and curve existence for a camera ray $\mathbf{r}$, which is given by the integrals
@@
$L_{\text{existence}}=$ $\frac{1}{N_{\text{rays}}}\sum_{\mathbf{r}}W_{\text{batch}}(\mathbf{r})||E(\mathbf{r})-\hat{E}(\mathbf{r})||_{2}^{2}$ (6)
@@
wher $E(\mathbf{r})$ and $\hat{E}(\mathbf{r})$ are the 2D-observed and predicted field existence, respectively, for both the surface and curve models, i indexes background pixels from input maps, j indexes the sample points along the rays, and $s$ determines the scale of the regularizer. The 2D-observed existence $E(\mathbf{r})$ is derived from the output of the 2D prior being a foreground pixel, either start, end, or barrel pixel.
@@
$\vspace{-0.3cm}W_{\text{batch}}(\mathbf{r})$ =$=\begin{cases}\frac{N_{\text{batch exist pixels}}}{N_{\text{batchpixels}}}&\text{ if }\mathbf{r}>\eta_{\text{batch}}\\ 1-\frac{N_{\text{batch exist pixels}}}{N_{\text{batch pixels}}}&\text{ otherwise}\end{cases}$ (7)
The weight $W_{\text{batch}}(r)$ addresses the imbalance between foreground and background areas in the existence maps, preventing the network from converging to a local minimum that falsely classifies most pixels as background, which are more prevalent.
Attribute Field $\mathcal{A}$
Now with the existence field, we can distill the information from the 2D priors by learning attribute fields $\mathcal{A}$ for both surfaces and curves, which we denote as $\mathcal{A}^{\text{surface}}$ and $\mathcal{A}^{\text{curve}}$. Following Zhi et al. [72], each attribute is modeled as an additional feature channel branching out from the existence field $\mathcal{F}$. For surfaces, the additional attribute field $\mathcal{A}^{\text{surface}}=\left\{p^{\text{surface}},q^{\text{surface}}\right\}$ is comprised of two networks $p^{\text{surface}}:\mathbb{R}^{3}\mapsto\mathbb{R}^{(K+1)}$ that predicts the extrusion instance label, and $q^{\text{surface}}:\mathbb{R}^{3}\mapsto\mathbb{R}^{4}$ that predicts the start-end-barrel segmentation. Similarly for the curves, the attribute field is given as $\mathcal{A}^{\text{curve}}=\left\{p^{\text{curve}},q^{\text{curve}}\right\}$, where $p^{\text{curve}}:\mathbb{R}^{3}\mapsto\mathbb{R}^{(K+1)}$ predicts the extrusion instance label, and $q^{\text{curve}}:\mathbb{R}^{3}\mapsto\mathbb{R}^{3}$ predicts the start-end segmentation. The expected attribute for a ray $\mathbf{r}$ is then computed using the volume rendering equation which is given as:
@@
for each $a\in\{p^{\text{surface}},q^{\text{surface}},p^{\text{curve}},q^{\text{curve}}\}$ that correspond to rendered predicted attributes $\hat{A}\in\{\bar{\mathbf{P}}^{\text{surface}},\bar{\mathbf{Q}}^{\text{surface}},\bar{\mathbf{P}}^{\text{curve}},\bar{\mathbf{Q}}^{\text{curve}}\}$ from the 3D field.
We train our attribute field using the predictions from our learned 2D priors $\mathcal{F}^{\text{surface}}$ and $\mathcal{F}^{\text{curve}}$. For each of the 2D multi-view images, we can extract the 2D observed labels $A\in\{\mathbf{P}^{\text{surface}},\mathbf{Q}^{\text{surface}},\mathbf{P}^{\text{curve}},\mathbf{Q}^{\text{curve}}\}$ from the learned priors, which we use as supervision. As mentioned in Sec. 3.2, since the labels may not individually be consistent across the multiple views, we use Hungarian matching during training to align the labels across multiple views based on the best one-to-one matching. For each attribute, we use a multi-class cross-entropy loss to train and optimize the attribute field with the given supervision:
@@
where $A^{l}(\mathbf{r})$ and $\hat{A}^{l}(\mathbf{r})$ represent the class $l$ attribute probabilities from the learned 2D prior and the predicted field, respectively, aligned with Hungarian matching.
@@
The integration into a coherent 3D field is optimized independently for each 3D shape. We use TensoRF [5] as the backbone representation for the 3D fields, where we have one model for the surface field and the other for the curve field. In both models, we train the existence fields $\mathcal{F}$ and the attribute fields $\mathcal{A}$ together for 1500 iterations. Each attribute field is parameterized by a 2-layer MLP that outputs logit values. Additional implementation details are provided in the appendix A5.2.
@@ 3.4 Reverse Engineering from 3D Fields
We now detail how we use our optimized surface field $\mathcal{F}^{\text{surface}}$ and curve field $\mathcal{F}^{\text{curve}}$ for the reverse engineering task, i.e., the recovery of the extrusion cylinders including the parameters $(\mathbf{n},\mathbf{c},\mathbf{h},\tilde{\mathbf{S}},\mathbf{s})$ of each. We first extract the surface point cloud and curve point cloud from the corresponding 3D fields by thresholding the corresponding existence fields. Each point in the point cloud would have attributes queried from the corresponding attribute fields. With these point clouds, the process to recover the extrusion cylinders is outlined as follows, with further details available in the appendix A5.3.
@@
    Extrusion Axis $\mathbf{n}$ Estimation: For each extrusion cylinder, the extrusion axis $\mathbf{n}$ is computed through plane fitting via RANSAC [12] on the surface base points with corresponding extrusion cylinder instance label. The axis $\mathbf{n}$ will then be set to fitted plane’s normal. This process relies on surface information because plane fitting tends to be more reliable with surface points. If the base face of the cylinder fails to reconstruct due to occlusion, we use curve points as a substitute, provided they exist.
@@
    2D Sketch $\tilde{\mathbf{S}}$ Estimation: To obtain the sketch $\tilde{\mathbf{S}}$ for each extrusion cylinder, we project the curve point cloud for each extrusion instance onto the plane with normal $\mathbf{n}$. From the projected point cloud, we can compute for the sketch scale $\mathbf{s}$. We then optimize an implicit signed distance function using the projected point cloud with IGR [15] per closed loop. Finally, we obtain the closed loop by extracting the zero-level set of the implicit function. The natural choice to extract the sketch is to use the curve information.
@@
    Extrusion height $\mathbf{h}$ Estimation: The height $\mathbf{h}$ of each extrusion cylinder is computed as the distance between the two sketch centers — from the start plane and the end plane. The sketch centers are computed from the projected curve point cloud for each paired labels (extrusion instance segmentation, start-end segmentation) onto the plane with normal $\mathbf{n}$. If any of the start-end planes are occluded or missing, the height is computed from the barrel points from the surface point cloud with the corresponding extrusion cylinder instance label.
@@ 4 Experiments
In this section, we present our experimental evaluation to demonstrate the capabilities of our MV2Cyl in reconstructing extrusion cylinders given only multi-view images. To the best of our knowledge, our proposed pipeline is the first to tackle extrusion cylinder recovery directly from 2D input images, and we show that our method is outperforms existing methods that take clean geometry as input and their modifications to handle raw input images.
@@
We evaluate our approach on two sketch-extrude CAD datasets Fusion360 [63] and DeepCAD [64]. We use the train and test splits as released in [58]. We enrich these datasets with multi-view image renderings and 2D curve and surface segmentation maps for the training and evaluation of our method and the baselines. We will release this dataset upon publication. More details about dataset generation is depicted in the appendix A5.4.
@@
Following [58], we report the extrusion-axis error (E.A.), extrusion-center error (E.C.), per-extrusion cylinder fitting loss (Fit Cyl.), and global fitting loss (Fit Glob.) for each shape. Additionally, we also report the extrusion-height error (E.H.) defined as the L1 distance between the GT and predicted height value: $\frac{1}{K}\sum_{k=1}^{K}|\hat{\mathbf{h}}_{k}-\mathbf{h}_{k}|$ while it is missed in Point2Cyl. These metrics evaluate how well the methods reconstruct and recover the parameters of the extrusion cylinders. The final scores are the average of the metrics calculated over all test shapes.
@@
  • NeuS2 [61]+Point2Cyl [58] combines NeuS2 [61] with Point2Cyl [58]. Since no prior work has reconstructed extrusion cylinders directly from multi-view images, we combine methods for shape reconstruction from multi-view images with techniques for extrusion cylinder reconstruction from raw geometry. This approach utilizes NeuS2 [61], an off-the-shelf image-based surface reconstruction technique, that optimizes the photometric loss to learn underlying implicit representations of 3D scenes. Specifically, it first reconstructs object surfaces from multi-view images and samples point clouds from the reconstructed mesh. The sampled points are directly fed into Point2Cyl [58] to predict sketches and extrusion parameters of CAD models.
  • Point2Cyl [58] is a technique for reconstructing sketch-extrude CAD given input 3D point clouds. We further compare the previous work of taking 3D explicit geometry(point cloud) as input. The clean point clouds sampled from the ground truth CAD meshes are utilized as input. It initially segments the input point cloud into distinct instances and base/barrel segments, then identifies the extrusion axis as the direction that aligns with the normals of the points in the base segments and is orthogonal to the normals of the points in the barrel segments. The rest of the parameters are inferred based on this predicted axis and the segmentation labels. To evaluate the network, we trained it using the official code released by the authors.
@@
Dataset Method E.A.$\downarrow$ E.C.$\downarrow$ E.H.$\downarrow$ Fit. Cyl. $\downarrow$ Fit. Glob.$\downarrow$

### v2

@@ Abstract
We present MV2Cyl, a novel method for reconstructing 3D from 2D multi-view images, not merely as a field or raw geometry but as a sketch-extrude CAD model. Extracting extrusion cylinders from raw 3D geometry has been extensively researched in computer vision, while the processing of 3D data through neural networks has remained a bottleneck. Since 3D scans are generally accompanied by multi-view images, leveraging 2D convolutional neural networks allows these images to be exploited as a rich source for extracting extrusion cylinder information. However, we observe that extracting only the surface information of the extrudes and utilizing it results in suboptimal outcomes due to the challenges in the occlusion and surface segmentation. By synergizing with the extracted base curve information, we achieve the optimal reconstruction result with the best accuracy in 2D sketch and extrude parameter estimation. Our experiments, comparing our method with previous work that takes a raw 3D point cloud as input, demonstrate the effectiveness of our approach by taking advantage of multi-view images. Our project page can be found at https://mv2cyl.github.io.
@@ 2 Related Work
Another line of work tackles fitting 3D parametric curves, where a common strategy is to first detect edges [70] and corners [35, 70, 39] from an input point cloud, and then fit or predict parametric curves from the initial proposals. A recent work, NEF [68], takes multi-view images as input and introduces a self-supervised approach to the parametric curve fitting problem using neural fields, thus allowing them to optimize without needing clean point cloud data. Despite having multi-view images as input, NEF only tackles recovering parametric curves and not extrusion cylinders, which is a basic building block for the sketch-extrude reverse engineering task and is the focus of our work.
@@ 3 MV2Cyl: Reconstructing 3D Extrusion Cylinders from Multi-View Images
Figure 1: Full Pipeline of MV2Cyl.
We propose a method that reconstructs 3D extrusion cylinders from multi-view images without relying on raw 3D geometry as input. The idea is to leverage 2D neural networks to learn 2D priors that provide 3D extrusion information (Sec. 3.2), i.e. extrusion curves and surfaces. The integration of the information into 3D is achieved through optimizing neural fields (Sec. 3.3). MV2Cyl is a combination of a curve field and a surface field that is used to recover the parameters and reconstruct the extrusion cylinders only given 2D input images (Sec. 3.4).
@@ 3.2 Learning 2D Priors for 3D Extrusions
We observe that detecting and extracting curves on the base planes of each extrusion cylinder segment provides additional information to recover the parameters and reconstruct the extrusion cylinders. Moreover, (feature) curves have been a longstanding and well-explored problem dating back to classical computer vision literature [46, 74, 43] thanks to their strong expressiveness. This leads us to learn a 2D curve prior that provides more discriminative and detailed outputs to make reverse engineering more robust.
@@
$\mathcal{L}_{\text{2D}}^{\text{curve}}=\lambda_{CE}\mathcal{L}_{\text{cross-entropy}}+\lambda_{\text{focal}}\mathcal{L}_{\text{focal}}+\lambda_{\text{dice}}\mathcal{L}_{\text{dice}}.\vspace{-0.1cm}$ (2)
Additional details about the segmentation framework are available in the appendix E.1.
@@ 3.3 Integrating 2D Segments into a 3D Field
Figure 3: Overview of the learned surface and curve fields. (Left-to-Right) Density field of surface, instance semantic field of surface, start-end semantic field of surface, density field of curve, instance semantic field of curve, and start-end semantic field of curve.
To establish correspondences and collate the information extracted from multi-view images, a natural approach to integrate 2D information into a consistent 3D representation is through learning a 3D field [40, 61, 68, 44]. For each of our 2D priors $\mathcal{M}^{\text{surface}}$ and $\mathcal{M}^{\text{curve}}$, we learn a density field $\mathcal{F}$ and a semantic field $\mathcal{A}$ as detailed below, respectively.
Density Field $\mathcal{F}$
We learn a density field for surfaces $\mathcal{F}^{\text{surface}}$ and curves $\mathcal{F}^{\text{curve}}$. Given a query point in 3D space $\mathbf{x}\in\mathbb{R}^{3}$, the density field network $\mathcal{F}:\mathbb{R}^{3}\mapsto\mathbb{R}$ outputs a scalar value that indicates how likely the 3D point $\mathbf{x}$ is on a surface or a curve for $\mathcal{F}^{\text{surface}}$ and curves $\mathcal{F}^{\text{curve}}$, respectively. To optimize the field differentiably with only 2D images, we use volume rendering [40, 5] and a learnable transformation [68] to convert the density field $\mathcal{F}(\mathbf{x})$ to an opacity field $\sigma(\mathbf{x})$. This transformation is in the form of a learnable sigmoid function given as:
$\vspace{-0.5pt}\sigma(\mathbf{x})=\alpha\cdot\left(1+e^{-\zeta\cdot(\mathcal{F}(\mathbf{x})-\beta)}\right)^{-1}\vspace{-0.5pt}$ (3)
where $\alpha$ is a learnable parameter that adjusts the density scale, and $\beta$ and $\zeta$ are constant hyperparameters. We use the same transformation in Eq. 3 for the surface field $\mathcal{F}^{\text{surface}}$ and the curve field $\mathcal{F}^{\text{curve}}$, resulting in corresponding opacity fields $\sigma^{\text{surface}}$ and $\sigma^{\text{curve}}$. We then directly render the density by using the volume rendering equation to get the expected surface and curve density for a camera ray $\mathbf{r}$, which is given by the integrals
@@
$L_{\text{density}}=$ $\frac{1}{N_{\text{rays}}}\sum_{\mathbf{r}}W_{\text{batch}}(\mathbf{r})||E(\mathbf{r})-\hat{E}(\mathbf{r})||_{2}^{2}$ (6)
@@
where $E(\mathbf{r})$ and $\hat{E}(\mathbf{r})$ are the 2D-observed and predicted field density, respectively, for both the surface and curve models, $i$ indexes background pixels from input maps, $j$ indexes the sample points along the rays, and $s$ determines the scale of the regularizer. The 2D-observed density $E(\mathbf{r})$ is derived from the output of the 2D prior being a foreground pixel, either start, end, or barrel pixel.
@@
$\vspace{-0.3cm}W_{\text{batch}}(\mathbf{r})$ =$=\begin{cases}\frac{N_{\text{batch exist pixels}}}{N_{\text{batchpixels}}}&\text{ if }\mathbf{r}>\eta_{\text{batch}}\\ 1-\frac{N_{\text{batch exist pixels}}}{N_{\text{batch pixels}}}&\text{ otherwise}\end{cases}\vspace{-0.1cm}$ (7)
The weight $W_{\text{batch}}(r)$ addresses the imbalance between foreground and background areas in the density maps, preventing the network from converging to a local minimum that falsely classifies most pixels as background, which are more prevalent.
Semantic Field $\mathcal{A}$
Now with the density field, we can distill the information from the 2D priors by learning semantic fields $\mathcal{A}$ for both surfaces and curves, which we denote as $\mathcal{A}^{\text{surface}}$ and $\mathcal{A}^{\text{curve}}$. Following Zhi et al. [73], each semantic is modeled as an additional feature channel branching out from the density field $\mathcal{F}$. For surfaces, the additional semantic field $\mathcal{A}^{\text{surface}}=\left\{p^{\text{surface}},q^{\text{surface}}\right\}$ is comprised of two networks $p^{\text{surface}}:\mathbb{R}^{3}\mapsto\mathbb{R}^{(K+1)}$ that predicts the extrusion instance label, and $q^{\text{surface}}:\mathbb{R}^{3}\mapsto\mathbb{R}^{4}$ that predicts the start-end-barrel segmentation. Similarly for the curves, the semantic field is given as $\mathcal{A}^{\text{curve}}=\left\{p^{\text{curve}},q^{\text{curve}}\right\}$, where $p^{\text{curve}}:\mathbb{R}^{3}\mapsto\mathbb{R}^{(K+1)}$ predicts the extrusion instance label, and $q^{\text{curve}}:\mathbb{R}^{3}\mapsto\mathbb{R}^{3}$ predicts the start-end segmentation. The expected semantic for a ray $\mathbf{r}$ is then computed using the volume rendering equation which is given as:
@@
for each $a\in\{p^{\text{surface}},q^{\text{surface}},p^{\text{curve}},q^{\text{curve}}\}$ that correspond to rendered predicted semantics $\hat{A}\in\{\bar{\mathbf{P}}^{\text{surface}},\bar{\mathbf{Q}}^{\text{surface}},\bar{\mathbf{P}}^{\text{curve}},\bar{\mathbf{Q}}^{\text{curve}}\}$ from the 3D field.
We train our semantic field using the predictions from our learned 2D priors $\mathcal{F}^{\text{surface}}$ and $\mathcal{F}^{\text{curve}}$. For each of the 2D multi-view images, we can extract the 2D observed labels $A\in\{\mathbf{P}^{\text{surface}},\mathbf{Q}^{\text{surface}},\mathbf{P}^{\text{curve}},\mathbf{Q}^{\text{curve}}\}$ from the learned priors, which we use as supervision. As mentioned in Sec. 3.2, since the labels may not individually be consistent across the multiple views, we use Hungarian matching during training to align the labels across multiple views based on the best one-to-one matching. For each semantic, we use a multi-class cross-entropy loss to train and optimize the semantic field with the given supervision:
@@
where $A^{l}(\mathbf{r})$ and $\hat{A}^{l}(\mathbf{r})$ represent the class $l$ semantic probabilities from the learned 2D prior and the predicted field, respectively, aligned with Hungarian matching.
@@
The integration into a coherent 3D field is optimized independently for each 3D shape. We use TensoRF [5] as the backbone representation for the 3D fields, where we have one model for the surface field and the other for the curve field. In both models, we train the density fields $\mathcal{F}$ and the semantic fields $\mathcal{A}$ together for 1500 iterations. Each semantic field is parameterized by a 2-layer MLP that outputs logit values. Additional implementation details are provided in the Appendix E.2.
@@ 3.4 Reverse Engineering from 3D Fields
We now elaborate on how we leverage our optimized surface field $\mathcal{F}^{\text{surface}}$ and curve field $\mathcal{F}^{\text{curve}}$ for the reverse engineering task, specifically to recover extrusion cylinders and their defining parameters, $(\mathbf{n},\mathbf{c},\mathbf{h},\tilde{\mathbf{S}},\mathbf{s})$. Each point within the extracted point clouds is assigned semantic information queried from the associated semantic fields, providing context for each feature in 3D space. Using these enriched point clouds, we employ a multi-step process to reconstruct each extrusion cylinder, iteratively refining the parameters to accurately fit the geometry and semantics. Further details can be found in Appendix E.3.
@@
    Extrusion Axis $\mathbf{n}$ Estimation: For each extrusion cylinder, the extrusion axis $\mathbf{n}$ is computed through plane fitting via RANSAC [12] on the surface base points with corresponding extrusion cylinder instance label. The axis $\mathbf{n}$ is then set as the normal of the fitted plane. This process relies on surface information because plane fitting is more reliable using the surface points. If the base face of the cylinder fails to be reconstructed due to occlusion, we use curve points as a substitute.
@@
    2D Sketch $\tilde{\mathbf{S}}$ Estimation: To obtain the sketch $\tilde{\mathbf{S}}$ for each extrusion cylinder, we first project the curve point cloud for each extrusion instance onto the plane with normal $\mathbf{n}$. From the projected point cloud, we then compute for the sketch scale $\mathbf{s}$. We then optimize an implicit signed distance function for each closed loop using the projected point cloud with IGR [15]. Finally, we obtain the closed loop by extracting the zero-level set of the implicit function. Implicit sketches can be converted to a parametric representation using an off-the-shelf spline fitting module. We provide example visualizations in Appendix  E.7.
@@
    Extrusion height $\mathbf{h}$ Estimation: The height $\mathbf{h}$ of each extrusion cylinder is computed as the distance between the two sketch centers — from the start plane and the end plane. The sketch centers are computed from the projected curve point cloud for each paired labels (extrusion instance segmentation, start-end segmentation) onto the plane with normal $\mathbf{n}$. If any of the start-end planes are occluded, the height is computed using the barrel points from the surface point cloud with the corresponding extrusion cylinder instance label.
@@
Figure 4: Converting 3D reconstructed geometry and semantics into CAD parameters.
@@ 4 Experiments
In this section, we present our experimental evaluation to demonstrate the feasibility of our MV2Cyl in reconstructing extrusion cylinders given only multi-view images. To the best of our knowledge, our proposed pipeline is the first to tackle extrusion cylinder recovery directly from 2D input images, and we show that our method is outperforms existing methods that take clean geometry as input and their modifications to handle raw input images.
@@
We evaluate our approach on two sketch-extrude CAD datasets Fusion360 [64] and DeepCAD [65]. We use the train and test splits as released in [58]. We enrich these datasets with multi-view image renderings and 2D curve and surface segmentation maps for the training and evaluation of our method and the baselines. More details about dataset generation are in Appendix E.4.
@@
Following [58], we report the average extrusion-axis error (E.A.), extrusion-center error (E.C.), per-extrusion cylinder fitting loss (Fit Cyl.), and global fitting loss (Fit Glob.). Additionally, we also report the extrusion-height error (E.H.) defined as the L1 distance between the GT and predicted height value: $\frac{1}{K}\sum_{k=1}^{K}|\hat{\mathbf{h}}_{k}-\mathbf{h}_{k}|$. These metrics evaluate how well the methods reconstruct and recover the parameters of the extrusion cylinders. The final scores are the average of the metrics calculated over all test shapes.
@@
  • We compare with Point2Cyl [58], a technique for reconstructing sketch-extrude CAD given input 3D point clouds. The point clouds are sampled from clean ground truth CAD meshes and are utilized as the input for this baseline. Point2Cyl [58] initially segments the input point cloud into distinct instances and base/barrel segments, then identifies the extrusion axis as the direction that aligns with the normals of the points in the base segments and is orthogonal to the normals of the points in the barrel segments. The rest of the parameters are inferred based on this predicted axis and the segmentation labels. To evaluate the network, we trained it using the official code released by the authors.
  • NeuS2 [62]+Point2Cyl [58] combines NeuS2 [62] with Point2Cyl [58]. Since no prior work has reconstructed extrusion cylinders directly from multi-view images, we combine methods for i) shape reconstruction from multi-view images, with techniques for ii) extrusion cylinder reconstruction from raw geometry. This approach utilizes NeuS2 [62], an off-the-shelf image-based surface reconstruction technique, that optimizes the photometric loss to learn underlying implicit 3D representations. For this baseline, we first reconstruct the object surface from multi-view images using NeuS2 [62], and then sample a point cloud from the reconstructed mesh. The sampled point cloud is then directly fed into Point2Cyl [58] to predict sketches and extrusion parameters of CAD models.
@@
Dataset Method E.A.($\downarrow$ E.C.$\downarrow$ E.H.$\downarrow$ Fit. Cyl. $\downarrow$ Fit. Glob.$\downarrow$